OpenStudy (anonymous):

Kim rode her bike at a rate of 24km/h to the post office and 16km/h back home. If her trip to the post office and back took a total of 75 minutes, how far does Kim live from the post office?

OpenStudy (anonymous):

Distance from home to post office: D Total time = D/24+D/16 = 5/4 5D/48=5/4 D=12 12km

OpenStudy (anonymous):

how did you u get 5/4

OpenStudy (anonymous):

75 minutes = 5/4 hours

OpenStudy (anonymous):

don't need to answer this anymore